OpenStudy (anonymous):

Cross Products and Sets?!HELPPP! ( I've done everything else perfectly fine but can't figure this out please help) Give an example of a situation where finding the cross product of two sets is useful and explain why it is useful. Your example should include two sets and their cross product.

OpenStudy (kitten_is_back):

okay i'll help you

OpenStudy (anonymous):

@Kitten_is_back I looked at that..and it wasn't what my instructor was looking for

OpenStudy (kitten_is_back):

okay

OpenStudy (kitten_is_back):

Give an example of a situation where finding the cross product of two sets is useful include two sets and there cross product ---- If an 8 ft tree casts a shadow of 20 ft, how long is the shadow of a 6 ft man? Use a proportion: x/6 = 20/8 x = 6(20/8) = 120/8 = 15 ft. =================

OpenStudy (anonymous):

Thank you! @Kitten_is_back

OpenStudy (kitten_is_back):

yw